# Stop! In the Name of Flaws: Disentangling Personal Names and Sociodemographic Attributes in NLP

### v1

@@ Abstract
Personal names simultaneously differentiate individuals and categorize them in ways that are important in a given society. While the natural language processing community has thus associated personal names with sociodemographic characteristics in a variety of tasks, researchers have engaged to varying degrees with the established methodological problems in doing so. To guide future work, we present an interdisciplinary background on names and naming. We then survey the issues inherent to associating names with sociodemographic attributes, covering problems of validity (e.g., systematic error, construct validity), as well as ethical concerns (e.g., harms, differential impact, cultural insensitivity). Finally, we provide guiding questions along with normative recommendations to avoid validity and ethical pitfalls when dealing with names and sociodemographic characteristics in natural language processing.
@@ 1 Introduction
A person’s identity is a complex and paradoxical thing - it simultaneously identifies someone’s uniqueness, and categorises them, identifying what they have in common with others (Strauss, 2017). A perfect example of this phenomenon is a person’s name. Personal names are proper nouns used to refer to individuals. They play an important distinguishing role in our lives, as they let us uniquely represent people mentally, refer to them directly in speech, and underscore their significance as individuals (Jeshion, 2009). For these reasons, personal names are a linguistic universal, i.e., they appear across languages and cultures, although naming customs vary across the world (Hough, 2016).
But alongside differentiating people, names also categorize them in their society. Names assigned to people often index aspects of identity that are important in the context of their society, including sex, religion, tribe, stage of life, etc. Personal names are thus rich resources to understand the social organization of communities, and have been studied across anthropology (Alford, 1987; Hough, 2016), sociology (Marx, 1999; Pilcher, 2017), linguistics (Allerton, 1987; Anderson, 2003) and onomastics (Alvarez-Altman et al., 1987; Adams, 2009).
@@
In natural language processing (NLP) as well, personal names have a long history of use—NLP researchers have worked on identifying and disambiguating uses of personal names (Mann and Yarowsky, 2003; Minkov et al., 2005; Färber and Ao, 2022) and have examined name translation (Sennrich et al., 2016; Wang et al., 2022; Sandoval et al., 2023) and name transliteration (Li et al., 2007). Increasingly, NLP researchers also use personal names along with sociodemographic characteristics for passive analysis of media and scholarly content (Vogel and Jurafsky, 2012; Knowles et al., 2016; Mohammad, 2020; Asr et al., 2021), or to examine model biases and harms (Maudslay et al., 2019; Romanov et al., 2019; Webster et al., 2021). However, these papers engage to varying degrees with concerns that have been raised outside of NLP about the methodological validity and ethics of associating names with sociodemographic characteristics. We argue that neglect of these issues is a significant barrier to valid and respectful research, as well as more inclusive NLP systems.
@@
We consider a range of harms in our work, from group-level, individual, and representational harms, to hermeneutical injustice and epistemic violence. We explain these terms where we use them in Section 4.
@@
All authors have a background in data science and ethics, and one has a background in philosophy. Three of the authors are trans and have names that are likely unintelligible to popular name-based sociodemographic inference methods. Two authors are trans people of color; as such, many examples in this paper reflect concerns about misgendering and racialization.
@@ 2.1 Names and Naming
Names are generally regarded as social phenomena that serve two central functions that are sometimes in conflict: differentiation and categorization of individuals (Alford, 1987). Differentiation is important psychologically and semantically for us to be able to directly refer to and mentally represent individuals, and names also serve to underscore their referent’s significance as an individual (Jeshion, 2009). Categorization, on the other hand, is important for the social organization of communities, and naming conventions tend to reflect factors that are important to a community at a given point in time, e.g., gender, religion, descent, transition to adulthood, and so on (Hough, 2016). Thus, names and naming can only be understood in a specific (geographic, cultural, temporal) context, and even then includes a lot of variation. As folk assumptions about names tend to overlook the wide variation in names and naming (McKenzie, 2010), we present an overview of naming as it relates to sociodemographic characteristics below.
@@
The markers considered important to index in a name vary widely across cultures. For example, almost all European naming systems and indeed most societies across the world tend to assign sex-typed names (Hough, 2016), while South Indian naming conventions often index caste (Meganathan, 2009). However, convention does not mean that every single individual is assigned a name that neatly follows that convention, as shown by the long history of gender-ambiguous names in the U.S. (Barry and Harper, 1982). Additionally, gendered associations for specific names change over time (Barry and Harper, 1993), as do naming conventions in societies–for example, it is becoming increasing popular to assign non-gendered names in the U.S. and in Israel (Hough, 2016; Obasi et al., 2019). Apart from conventions, names themselves are not static and unchanging from birth, with many names changing due to partnerships, adoption, transition to a different life stage or gender, and so on (AIATSIS, 2022; Obasi et al., 2019; Hough, 2016).
@@
Trends in big-picture naming conventions are complicated by factions of society who want to resist imposed classification. Increasingly heterogeneous societies are a natural setting for such tensions; cross-cultural associations with sociodemographic characteristics can differ and sometimes clash, complicating naming, e.g., names like Nicola and Andrea tend to be assigned to boys in Italy but to girls in Germany. As Germany is a society with highly regulated naming practices, inclusion of these names necessitated a court judgment Hough (2016). Immigrant families thus have to juggle the delicate balance of asserting their identity but avoiding name-based stigma and discrimination in the new culture. Their naming practices have therefore been studied as an indicator of attitudes towards assimilation or its rejection, showing how names are not a transparent indicator of race (Sue and Telles, 2007; Becker, 2009). Even among adults, imperialism and colonialism are forces that affect naming. Indigenous individuals have been forced to adopt Western names in settler colonial and post colonial societies, e.g., US, Canada, Australia (AIATSIS, 2022). Similarly, Chinese individuals around the world adopt Western names in conversational (Li, 1997) and professional settings (Chan, 2016). Among trans and gender-nonconforming adults, many choose a new name to reflect and express their gender, walking the tightrope between normativity and self-assertion (Konnelly, 2021); Obasi et al. (2019) find that 50% of gender-nonconforming respondents who change their name pick a gender-neutral name. Beyond transgender people, new names and pseudonyms are also often self-selected to assert agency in identity creation, e.g., bell hooks, Sojourner Truth, and Malcolm X (Baker and Green, 2021).
@@
As naming involves a trade-off between differentiation and categorization, names often recur, a quantitative assumption that a lot of sociological, anthropological and NLP classification relies on (Alford, 1987). However, the distributions of names and people are very different. Weitman (1981) finds that in 100 years of first names from Israel’s Population Registry, the most frequent names (101+ occurrences) account for the majority of the population of a society (91%), but this corresponds to just a tiny minority of all assigned names (2.93%). These numbers could vary widely depending on the society, as, for example, the Chuukese people of Micronesia have a tradition of giving entirely unique names to children (Alford, 1987). Hence, it is important to distinguish when names are the object of study and when people are, to contextualize any results that involve the analysis of names.
@@ 2.2 Names and Sociodemographic Characteristics in NLP
Here, we present a non-comprehensive list of papers to illustrate some common uses of names and sociodemographic characteristics in NLP. We take an expansive view of “NLP,” including instances of machine processing of natural language without limiting ourselves to any particular venues.
@@
Numerous NLP works have developed algorithms to infer sociodemographic attributes from names (Chang et al., 2010; Liu and Ruths, 2013; Knowles et al., 2016), e.g., for passive analysis of social media content. Another line of NLP papers have relied on names to quantify gender disparities in academic publishing (Vogel and Jurafsky, 2012; Mohammad, 2020) or media representation (Asr et al., 2021). Some NLP works have identified preserving hegemonic gender associations as an important criterion for transliteration and translation (Li et al., 2007; Wang et al., 2022). Furthermore, names are increasingly leveraged to investigate social biases in NLP models. For example, De-Arteaga et al. (2019) study how first names, which they consider “explicit gender indicators,” affect the gender bias of occupation prediction from biographies. Similarly, Jeoung et al. (2023) assess the causal impact of first names, which they posit “may serve as proxies for (intersectional) socio-demographic representations,” on the commonsense reasoning performance of language models. Smith and Williams (2021) measure racial biases as well, evaluating generative dialogue models by having “one conversational partner […] state a name commonly associated with a certain gender and/or race/ethnicity.” In this line of research, it is commonplace to use skewed reference populations such as U.S. census data (U.S. Census, 2020) and Social Security Administration baby names (U.S. Social Security Administration, 2023) for gender assocations (Lockhart et al., 2023).
@@
In these works, researchers have engaged to varying degrees with the established methodological and ethical problems of associating names with sociodemographic characteristics. Some papers have investigated biases in algorithms that aim to infer sociodemographic attributes from names; for example, Kozlowski et al. (2022) measure racial bias in inferring race from names in scientific articles, finding Black authors are underestimated while white authors are overestimated. Knowles et al. (2016) state that their classifier to predict gender from names is biased towards the U.S. and assumes gender is binary, but leaves these issues “to be addressed in future work.” Mohammad (2020) acknowledge that inferring gender from names can yield misgendering because “names do not capture gender fluidity or contextual gender,” but suggest a tradeoff with “the benefits of NLP techniques and social category detection.” Despite this critical reflection, NLP papers also make unfounded assumptions about names. For instance, Vogel and Jurafsky (2012) posit that certain names are “unambiguous” with respect to gender across languages, and Wang et al. (2022) claim that there exist “names with obvious gender.”
3 Issues of Validity
@@
Error is not robustly quantifiable.
Due to the wide range of variation in naming culturally, temporally, and geographically, mislabelling happens at different rates for different individuals and groups (c.f., random errors). Multiple studies empirically analyze the error rates of name-based gender and race inference systems as compared to gold data [^1] All these studies look at imputing an individual’s gender, but the gold labels they compare to are, confusingly, not always self-reported gender! Some use gender assigned by annotators as the ground truth, which would be fine if comparing to perceptions of an individual based on their name, but these studies do not, raising further questions about their methodological validity. in different contexts (Karimi et al., 2016; Kozlowski et al., 2022; Van Buskirk et al., 2023; Lockhart et al., 2023). Lockhart et al. (2023) find that depending on the tool, overall error rates range from 4.6% to 86% overall, and up to 100% for particular subgroups. This extremely wide range of error shows how unreliable names are in general as a method, and also highlights how contextual the error rates are, i.e., a manual analysis with self-reported data would be needed to quantify the error rate of a system using names in a new environment. Otherwise, it is simply not possible to reasonably estimate the bounds of error of a name-based analysis, and results with it should not be taken seriously.
Systematic error and selection bias.
Many names are simply not informative of certain sociodemographic variables, in given (social, geographical and temporal) contexts and with homogeneous populations. In the context of gender, this includes neutral names like Alex with no unique gendered association in the U.S. and Canada, and with race, there is overlap between the names Black and white parents give to children in the U.S. (Lockhart et al., 2023), and religious names are used in many parts of the world (Curtis, 2005; Olúwáfeḿi, 2014). At the intersection of gender and race, many Chinese names when Romanized are not gender-associated, and infrequent names (i.e., the vast majority of names!), are also not informative. These “edge cases” are sometimes addressed by just assigning the majority class label; if the name Miaoran is “60% female” according to a gender prediction tool, some NLP systems use this to classify all people named Miaoran as female, resulting in systematic error (Kirkup and Frenkel, 2006). Alternatively, uninformative names are simply excluded from the analysis, resulting in selection bias. This completely alters the makeup of the data and therefore the results (Mihaljević et al., 2019). Both these preprocessing choices are examples of data cleaning procedures that affect internal validity, i.e., gaps in the translation from measurements to overall conclusions (Liao et al., 2021), leading to less robust and trustworthy results.
@@
Construct validity asks how well an abstract concept can be measured through some indicator (Messick, 1995). In our case, it is obvious to ask how valid it is to assign sociodemographic categories via names (in a given context), but construct validity also applies to the sociodemographic categories themselves (Benthall and Haynes, 2019; Hanna et al., 2020), which often reduce multiple dimensions of interest to a single label, e.g., “gender” might conflate identity, socialization, expression and perception (Keyes et al., 2021), which could all potentially differ. For instance, Mohammad (2020) uses names to index both identity (to investigate trends in authorship) and perception (to investigate trends in citation), which need not be the same, in a bibliometric analysis of the ACL Anthology. Similarly, a person might experience different racialized harms based on their perceived race, compared to their racial identity. As identity categories do not always line up neatly with experiences of marginalization, different dimensions might be relevant depending on the research question, and the goal. Another infrequently considered aspect of construct validity is the validity of abstract concepts such as “bias” and “fairness,” which show up frequently in the study of names and sociodemographic categories in NLP (Blodgett et al., 2020; Jacobs and Wallach, 2021).

### v2

@@ Abstract
Personal names simultaneously differentiate individuals and categorize them in ways that are important in a given society. While the natural language processing community has thus associated personal names with sociodemographic characteristics in a variety of tasks, researchers have engaged to varying degrees with the established methodological problems in doing so. To guide future work that uses names and sociodemographic characteristics, we provide an overview of relevant research: first, we present an interdisciplinary background on names and naming. We then survey the issues inherent to associating names with sociodemographic attributes, covering problems of validity (e.g., systematic error, construct validity), as well as ethical concerns (e.g., harms, differential impact, cultural insensitivity). Finally, we provide guiding questions along with normative recommendations to avoid validity and ethical pitfalls when dealing with names and sociodemographic characteristics in natural language processing.
@@ 1 Introduction
A person’s identity is a complex and paradoxical thing - it simultaneously identifies someone’s uniqueness, and categorizes them, identifying what they have in common with others (Strauss, 2017). A perfect example of this phenomenon is a person’s name. Personal names are proper nouns used to refer to individuals. They play an important distinguishing role in our lives, as they let us uniquely represent people mentally, refer to them directly in speech, and underscore their significance as individuals (Jeshion, 2009). For these reasons, personal names are a linguistic universal, i.e., they appear across languages and cultures, although naming customs vary across the world (Hough, 2016).
But alongside differentiating people, names also categorize them in their society. Names assigned to people often index aspects of identity that are important in the context of their society, including sex, religion, tribe, stage of life, etc. Personal names are thus rich resources to understand the social organization of communities, and have been studied across anthropology (Alford, 1987; Hough, 2016), sociology (Marx, 1999; Pilcher, 2017), linguistics (Allerton, 1987; Anderson, 2003), and onomastics (Alvarez-Altman et al., 1987; Adams, 2009).
@@
In natural language processing (NLP) as well, personal names have a long history of use—NLP researchers have worked on identifying and disambiguating uses of personal names (Mann and Yarowsky, 2003; Minkov et al., 2005; Färber and Ao, 2022) and have examined name translation (Sennrich et al., 2016; Wang et al., 2022; Sandoval et al., 2023) and name transliteration (Li et al., 2007; Benites et al., 2020; Sälevä and Lignos, 2024). Increasingly, NLP researchers also use personal names along with sociodemographic characteristics for passive analysis of media and scholarly content (Vogel and Jurafsky, 2012; Knowles et al., 2016; Mohammad, 2020; Asr et al., 2021), or to examine model biases and harms (Maudslay et al., 2019; Romanov et al., 2019; Webster et al., 2021). However, these papers engage to varying degrees with concerns that have been raised outside of NLP about the methodological validity and ethics of associating names with sociodemographic characteristics. We argue that neglect of these issues is a significant barrier to valid and respectful research, as well as more inclusive NLP systems.
@@
We consider group-level, individual, and representational harms, terms which we explain where we use them in Section 4.
@@
All authors have a background in data science and ethics, and one has a background in philosophy. Three of the authors are trans and have names that are likely unintelligible to popular name-based sociodemographic inference methods. Two authors are trans people of colour; as such, many examples in this paper reflect concerns about misgendering and racialization.
@@ 2.1 Names and Naming
Names are generally regarded as social phenomena that serve two central functions that are sometimes in conflict: differentiation and categorization of individuals (Alford, 1987). Differentiation is important psychologically and semantically for us to be able to directly refer to and mentally represent individuals, and names also serve to underscore their referent’s significance as an individual (Jeshion, 2009). Categorization, on the other hand, is important for the social organization of communities, and naming conventions tend to reflect factors that are important to a community at a given point in time, e.g., gender, religion, descent, transition to adulthood, and so on (Hough, 2016). For instance, the practice of naming someone after their father or grandfather—patronymic naming—was once common across Europe, and was popular in Sweden until the nineteenth century (e.g., Samuelsson) and continues into Iceland today (e.g., Gunnarsdóttir(Hough, 2016). This example shows how names and naming can only be understood in a specific (geographic, cultural, temporal) context, and even then includes a lot of variation. As folk assumptions about names tend to overlook the wide variation in names and naming (McKenzie, 2010), we present an overview of naming as it relates to sociodemographic characteristics below.
@@
The markers considered important to index in a name vary widely across cultures. For example, almost all European naming systems and indeed most societies across the world tend to assign sex-typed names (Hough, 2016), while South Indian naming conventions often index caste (Meganathan, 2009). However, convention does not mean that every single individual is assigned a name that neatly follows that convention, as shown by the long history of gender-ambiguous names in the U.S. (Barry and Harper, 1982). Additionally, gendered associations for specific names change over time (Barry and Harper, 1993), as do naming conventions in societies–for example, it is becoming increasingly popular to assign non-gendered names in the U.S. and in Israel (Hough, 2016; Obasi et al., 2019). Apart from conventions, names themselves are not static and unchanging from birth, with many names changing due to partnerships, adoption, transition to a different life stage or gender, and so on (Hough, 2016; Obasi et al., 2019; AIATSIS, 2022).
@@
Trends in big-picture naming conventions are complicated by factions of society who want to resist imposed classification. Increasingly heterogeneous societies are a natural setting for such tensions; cross-cultural associations with sociodemographic characteristics can differ and sometimes clash, complicating naming, e.g., names like Nicola and Andrea tend to be assigned to boys in Italy but to girls in Germany. As Germany is a society with highly regulated naming practices, inclusion of these names necessitated a court judgment Hough (2016). Immigrant families thus have to juggle the delicate balance of asserting their identity but avoiding name-based stigma and discrimination in the new culture. Their naming practices have therefore been studied as an indicator of attitudes towards assimilation or its rejection, showing how names are not a transparent indicator of race (Sue and Telles, 2007; Becker, 2009). Even among adults, imperialism and colonialism are forces that affect naming. Indigenous individuals have been forced to adopt Western names in settler colonial and postcolonial societies, e.g., the U.S., Canada, Australia (AIATSIS, 2022). Similarly, Chinese individuals around the world adopt Western names in conversational (Li, 1997) and professional settings (Chan, 2016). Among trans and gender-nonconforming adults, many choose a new name to reflect and express their gender, walking the tightrope between normativity and self-assertion (Konnelly, 2021); Obasi et al. (2019) find that 50% of gender-nonconforming respondents who change their name pick a gender-neutral name. Beyond transgender people, new names and pseudonyms are also often self-selected to assert agency in identity creation, e.g., bell hooks, Sojourner Truth, and Malcolm X (Baker and Green, 2021).
@@
As naming involves a trade-off between differentiation and categorization, names often recur, a quantitative assumption that a lot of sociological, anthropological and NLP classification relies on (Alford, 1987). However, the distributions of names and people can be very different. Weitman (1981) finds that in 100 years of first names from Israel’s Population Registry, the most frequent names (101+ occurrences) account for the majority of the population of a society (91%), but this corresponds to just a tiny minority of all assigned names (2.93%). These numbers could vary widely depending on the society, as, for example, the Chuukese people of Micronesia have a tradition of giving entirely unique names to children (Alford, 1987). Hence, it is important to distinguish when names are the object of study and when people are, to contextualize any results that involve the analysis of names.
@@ 2.2 Names and Sociodemographic Characteristics in NLP
Here, we present a non-comprehensive list of papers to illustrate some common uses of names and sociodemographic characteristics in NLP.
@@
Numerous NLP works have developed algorithms to infer sociodemographic attributes from names (Chang et al., 2010; Liu and Ruths, 2013; Knowles et al., 2016), e.g., for passive analysis of social media content. Another line of NLP papers have relied on names to quantify gender disparities in academic publishing (Vogel and Jurafsky, 2012; Mohammad, 2020) or media representation (Asr et al., 2021). Some NLP works have identified preserving dominant gender associations as an important criterion for transliteration and translation (Li et al., 2007; Wang et al., 2022). Names are also used to investigate social biases in NLP systems and language models (Kotek et al., 2023; An et al., 2023; Ibaraki et al., 2024). For example, De-Arteaga et al. (2019) study how first names, which they consider “explicit gender indicators,” affect the gender bias of occupation prediction from biographies. Similarly, Jeoung et al. (2023) assess the causal impact of first names, which they posit “may serve as proxies for (intersectional) socio-demographic representations,” on the commonsense reasoning performance of language models. Smith and Williams (2021) measure racial biases as well, evaluating generative dialogue models by having “one conversational partner […] state a name commonly associated with a certain gender and/or race/ethnicity.” In this line of research, it is commonplace to use skewed reference populations such as U.S. census data (U.S. Census, 2020) and Social Security Administration baby names (U.S. Social Security Administration, 2023) for gender assocations (Lockhart et al., 2023).
@@
In these works, researchers engage to varying degrees with the established methodological and ethical problems of associating names with sociodemographic characteristics. Some NLP papers make unfounded assumptions about names, e.g., Vogel and Jurafsky (2012) posit that certain names are “unambiguous” with respect to gender across languages, and Wang et al. (2022) claim that there exist “names with obvious gender.” Other papers are more critically reflective, acknowledging the limitations of their work: Knowles et al. (2016) state that their classifier to predict gender from names is biased towards the U.S. and assumes gender is binary, but leaves these issues “to be addressed in future work.” Mohammad (2020) acknowledges that inferring gender from names can yield misgendering because “names do not capture gender fluidity or contextual gender,” but suggest a trade-off with “the benefits of NLP techniques and social category detection.” Encouragingly, some recent papers opt for more inclusive study designs after engaging deeply with the pitfalls of using names and sociodemographic characteristics (Sandoval et al., 2023; Saunders and Olsen, 2023; Lassen et al., 2023).
3 Validity Issues
@@
Error is not quantifiable without asking humans.
The accuracy of using names to infer sociodemographic characteristics of individuals cannot be quantified without ground truth data, which for people’s identities, can only be obtained by asking them. Multiple studies thus empirically analyze the error rates of name-based gender and race inference systems as compared to gold data in different contexts (Karimi et al., 2016; Kozlowski et al., 2022; Van Buskirk et al., 2023; Lockhart et al., 2023). [^1] All these studies look at imputing an individual’s gender, but the gold labels they compare to are, confusingly, not always self-reported gender! Some use gender assigned by annotators as the ground truth, which would be fine if comparing to perceptions of an individual based on their name, but these studies do not, raising further questions about their methodological validity. For example, Lockhart et al. (2023) evaluate gender and race inference systems using self-reported data from nearly 20,000 individuals. Importantly, their self-reported data does not directly transfer to other contexts, as their respondents are authors of English language social science journal articles who are mostly located in the U.S. Using this data as reference data for a system with users located primarily in India, or for U.S. authors in a different century, makes little sense. In new environments, it is simply not possible to reasonably estimate the bounds of error of a name-based analysis, and results without a corresponding analysis of self-reported data should not be taken seriously.
Popular design choices lead to systematic error and selection bias.
Names that are uninformative of a sociodemographic characteristic present an issue for tools that aim to label everyone. In the context of gender, names like Alex have no unique gendered association in the U.S. and Canada; with race, names assigned by Black and white parents overlap in the U.S. (Lockhart et al., 2023), and religious names are used around the world (Curtis, 2005; Olúwáfeḿi, 2014)); at the intersection of gender and race, many Chinese names are not gender-associated when Romanized, and infrequent names are also not informative. Two common design choices for handling uninformative names are to assign the majority class label anyway, or, alternatively, to just exclude them. Assigning the majority class (i.e., classifying all people named Miaoran as female if a gender prediction tool predicts the name to be “60% female”) results in systematic error (Kirkup and Frenkel, 2006). On the other hand, excluding uninformative names from the analysis completely alters the makeup of the data and therefore the results (Mihaljević et al., 2019), resulting in selection bias. Both choices affect internal validity, i.e., gaps in the translation from measurements to overall conclusions (Liao et al., 2021), leading to less robust and trustworthy results.
@@
Construct validity asks how well an abstract concept can be measured through some indicator (Messick, 1995); in our case, the question is: how valid is it to assign sociodemographic categories via names? [^2] While we focus on the construct validity of names in this section, we note that poor construct validity also applies to the sociodemographic categories themselves (Benthall and Haynes, 2019; Hanna et al., 2020) and to abstract concepts such as “bias” and “fairness,” which show up frequently in the study of names and sociodemographic categories in NLP (Blodgett et al., 2020; Jacobs and Wallach, 2021). The answer to this depends on what aspects of the sociodemographic category we are interested in: identity, socialization, expression, perception—all of which could differ and are frequently conflated (Keyes et al., 2021). As discussed previously, many names are simply not informative of certain sociodemographic identities in given contexts and with homogeneous populations; Lockhart et al. (2023) find that overall error rates of name-based gender and race imputation tools range from 4.6% to 86% overall, and up to 100% for particular subgroups, depending on the tool. However, when it comes to the perception of names as indexing a sociodemographic category, some names may have stronger construct validity, an assumption used by Sandoval et al. (2023) in their examination of names assigned at birth that are strongly associated with the baby’s sex and the parents’ race/ethnicity. On the other hand, Mohammad (2020) uses names to operationalize both identity (to investigate trends in authorship) and perception (to investigate trends in citation) in a bibliometric analysis of the ACL Anthology, even though these need not match, and many underrepresented names are uninformative of identity as well as perception (Van Buskirk et al., 2023). As names do not neatly line up with sociodemographic identities, perceptions, or experiences in a context-independent way, it is critical to investigate construct validity of names in any setting where they are used.